# A Moment-Conserving Discontinuous Galerkin Representation of the Relativistic Maxwellian Distribution

### v1

@@ 4.3 Relativistic BGK
Choosing a constant collision frequency of $\nu=1$, we can exactly solve equation 30, $f(t,u)=f^{MJ}(u)+exp(-\nu t)\left(f^{WB}(u)-f^{MJ}(u)\right)$. The solution can be interpreted as the difference between the initial water-bag distribution and the Maxwell-Jüttner distribution, which vanishes the discrepancy with time exponentially until only the Maxwell-Jüttner distribution remains. Figure 4 shows the evolution of $f(u)$ at three separate points that match the exact solution of equation 30 over time.
@@ Acknowledgements
This work was supported by the U.S. Department of Energy under contract number DE-AC02-09CH11466. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. G.J. was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021110.

### v2

@@ 4.3 Relativistic BGK
Choosing a constant collision frequency of $\nu=1$, we can exactly solve equation (30), $f(t,u)=f^{MJ}(u)+exp(-\nu t)\left(f^{WB}(u)-f^{MJ}(u)\right)$. The solution can be interpreted as the difference between the initial water-bag distribution and the Maxwell-Jüttner distribution, which vanishes the discrepancy with time exponentially until only the Maxwell-Jüttner distribution remains. Figure 4 shows the evolution of $f(u)$ at three separate points that match the exact solution of equation (30) over time.
@@ 5 Conclusion
The correction routine further makes possible simulations with reasonable domains and in multiple momentum space dimensions by reducing the grid requirements per dimensions needed to maintain machine precision accurate moments. For a distribution without corrected velocity and temperature moments, Figure 1 illustrated convergence to machine precision matched moments takes 10 000 cells and domains of $u_{max}=\pm 160$. While Figure 4 demonstrated that with the correction routine, a domain of only $u_{max}=\pm 4$ and 32 cells could maintain machine precision in the moments over many iteration of the projection routine.
@@ Acknowledgements
The authors thank the Gkeyll team, especially Dingyun Liu and Greg Hammett, for useful discussions regarding correction routines.
Funding
This work was supported by the U.S. Department of Energy under contract number DE-AC02-09CH11466. G.J. was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0021110.
Declaration of interests
The authors report no conflict of interest.